# Trajectory First: A Curriculum for Discovering Diverse Policies

### v1

@@ Abstract
Being able to solve a task in diverse ways makes agents more robust to task variations and less prone to local optima. In this context, constrained diversity optimization has emerged as a powerful reinforcement learning (RL) framework to train a diverse set of agents in parallel. However, existing constrained-diversity RL methods often under-explore in complex tasks such as robotic manipulation, leading to a lack in policy diversity. To improve diversity optimization in RL, we therefore propose a curriculum that first explores at the trajectory level before learning step-based policies. In our empirical evaluation, we provide novel insights into the shortcoming of skill-based diversity optimization, and demonstrate empirically that our curriculum improves the diversity of the learned skills.
@@ 1 Introduction
Reinforcement Learning (RL) has driven breakthroughs in robot locomotion [1], game-playing [2, 3], and foundation-model finetuning [4]. While effective, most RL methods assume a unimodal action distribution and produce only a single policy. In contrast, humans and animals routinely solve the same task using multiple qualitatively different strategies. Such variability is also desirable in RL, as it has been shown that human strategy diversity increases solution quality and robustness [5, 6]. Therefore, this work considers the discovery of a policy set that maximizes the reward in diverse ways.
@@
A number of previous works have investigated this problem from various perspectives. Notably, the fields of Novelty Search (NS) and Quality-Diversity (QD) have proposed a multitude of algorithms which populate an archive of solutions based on their novelty and performance [7, 8, 9, 10]. Further, gradient-based RL approaches define intrinsic diversity rewards that they combine with extrinsic task rewards using Lagrange multipliers [11], bandits [12], or linear combinations [13, 14, 15].
While effective, the above approaches are not without shortcomings. QD may produce exceptional results, but often at the cost of sample efficiency and manual feature design. Gradient-based diversity or entropy bonuses in RL may still collapse to a few modes and remain under-evaluated in challenging contact-rich tasks [16]. Since skill learning and diversity or entropy maximization are common exploration objectives (see Zheng et al. [17] for an overview), this observation may appear counterintuitive at first. Yet, we corroborate these findings in our work, demonstrating that even simple obstacle avoidance tasks may be substantially challenging for diversity optimization in RL (Fig. 2). Further, we highlight the necessity for exploration strategies tailored for diversity optimization, since we find that out-of-the-box exploration methods do well in improving task performance, but do not guarantee policy diversity.
@@
To overcome these shortcomings, we propose a new and simple curriculum for learning diverse policies with off-policy and model-free RL (Section 5). We motivate our approach by the observation that skill-based RL methods commonly explore by optimizing neural-network parameters which can be inefficient. In contrast, we introduce a trajectory-first curriculum that (i) uses an evolutionary search over open-loop action sequences to uncover a diverse set of high-reward behaviors, and (ii) distills these behaviors into distinct, off-policy, model-free policies. While prior work proposed similar formulations that first solve exploration and then learning [18, 19], we do not rely on human demonstrations and propose an evolutionary approach to maximize entropy by optimizing trajectory parameters for novelty instead of neural network parameters. Based on this algorithm, we empirically highlight shortcomings of existing diversity optimization methods in Section 5 and illustrate how the proposed curriculum enables learning diverse sets of robot manipulation policies.
@@
In short, we make three contributions with this work. First, we propose a novel curriculum for diversity optimization under extrinsic task rewards (see Fig. 1). Second, we introduce Constrained Novelty Search (CNS) to discover diverse trajectories in the first step of this curriculum (Section 3.1). Finally, we investigate how diversity can be maintained during online training of control policies from this data (Section 3.2).
@@
Figure 1: Overview of the proposed diversity curriculum. First, an evolution strategy explores in trajectory space to maximize novelty of trajectory parameters $\omega\in\Omega$ under performance constraints. Then, this data is used to warmstart the online training of multiple RL agents $\pi\in\Pi$ to solve the same optimization problem in policy space.
@@ 2 Preliminaries
While earlier works used scalars to balance diversity and task rewards, Zahavy et al. [11] introduced the following constrained MDP formulation:
@@
where $\Pi^{n}$ is the current set of policies, $v^{*}\triangleq\max_{\pi\in\Pi}J(\pi)$ is the value of the optimal policy and $\alpha\in[0,1)$ is a hyperparameter controlling the optimality constraint. This constrained optimization problem can then be solved efficiently using Lagrange multipliers that are tuned using dual ascent [22, 23]. Similar to subsequent work [17], we adopt the same problem formulation in this work. To quantify diversity, we will measure the distance to the nearest neighbor, which shall be maximized:
@@
Most novelty-based approaches to skill learning maximize the entropy of the policy set by using the entropy of the current policy set as an intrinsic reward [10, 8, 27]. To quantify entropy, the particle-based entropy estimator by Singh et al. [28] is commonly employed which estimates the sparsity of the distribution based on the distance between the datapoints $\{x_{i}\}_{i=1}^{n}$ and their $k$-nearest neighbor: ${\cal H}_{\text{particle}}(X)\propto\sum_{x_{i}\in X}\log\lVert x_{i}-x_{i}^{(k)}\rVert$. In practice, we choose $k=1$. In other words, we measure the particle-based entropy as distance to the nearest neighbor, and add a constant $c=1$ for numerical stability:
@@ 3 Evolutionary Exploration for Diverse Policy Discovery
Solving the problem in Eq. 1 requires an initialization that is sufficiently diverse to prevent the diversity optimization from only occurring locally. Since the policies are initialized randomly at the beginning, we find that the discovered policy set rarely covers the task space sufficiently and instead focuses on a subset of greedy solutions. While shaping the extrinsic reward is an option to encourage diverse interactions between agent and environment, tuning such reward functions is a tedious task. We propose an alternative approach to this, which $(i)$ uses an evolution strategy (ES) to explore in the space of trajectories before $(ii)$ learning a diverse set of policies from this data. We provide an intuition for this approach in Fig. 1 and describe both stages of our curriculum in the following. For a formal algorithmic description of the method, we refer to Appendix A.
@@ 3.1 Constrained Novelty Search for Spline-based Exploration
The first stage of our curriculum directly optimizes agent trajectories $\tau\in{\mathbb{R}}^{T\times u}$, where $T$ denotes the number of timesteps and $u$ the robot’s degrees of freedom. We represent a trajectory as a B-spline [29] parameterized by a control point matrix $\omega\in{\mathbb{R}}^{m\times u}$. Following the constrained diversity optimization objective in Eq. 1, we optimize a set of trajectory parameters $\Omega=\{\omega_{i}\}_{i=1}^{n}$ such that the resulting trajectories $\tau(\omega)$ are as diverse as possible under the constraint of near-optimality:
@@
where $J(\tau(\omega_{i}))=\textstyle\sum_{t}r_{\text{ext}}(s^{i}_{t},a^{i}_{t})$, and $v^{*}=\max_{\omega}J(\tau(\omega))$. Since the extrinsic task reward function $r_{\text{ext}}$ is generally non-differentiable we optimize this problem using a multi-population evolution strategy (ES). Following Zahavy et al. [11] we solve the dual optimization using gradient ascent with bounded Lagrange multipliers $\{\lambda_{i}\}_{i=1}^{n}$, which yields the following trajectory reward that we evaluate on the trajectory $\tau_{i}=\tau(\omega_{i})$:
@@
where $\phi$ is a feature extraction function that projects the states to a lower dimension, $\lambda_{i}$ is the $i$-th Lagrange multiplier, and $\sigma$ denotes the sigmoid function $\sigma(x)=1/(1+\exp(-x))$. We note that Eq. 7 is a generalized version of the novelty search objective from Conti et al. [10], but using population-level Lagrange multipliers instead of a single heuristically selected scalar. Using Lagrange multipliers permits not only to consider different weights for different populations, but also a dynamic adaptation of these weightings. We denote this objective and its optimization as Constrained Novelty Search (CNS) in the following. While Zahavy et al. [11] derive the analytical gradient of Eq. 2 for their intrinsic reward, CNS approximates this gradient by stochastic sampling from an ES, which approximates natural gradient steps on the novelty objective [30, 31, 32].
@@
Finally, we note again that we optimize in ${\mathbb{R}}^{m\times u}$ instead of ${\mathbb{R}}^{d}$ at this step where $d$ is the dimension of policy parameters. Since typically $m\times u\ll d$, we can optimize Eq. 7 with the CMA-ES [32], which is empirically more sample efficient in moderately high dimensional parameter spaces than using isotropic search distributions [35, 10] .
@@ 3.2 Efficient Online Diversity Optimization from Prior Data
Given a diverse dataset ${\cal D}=\{(\tau_{i},\,z_{i},\,r_{\mathrm{ext}}(\tau_{i}))\}_{i=1}^{\ell}$ with skill labels $z_{i}\in\{1,\dots,n\}$ from CNS, our second stage learns $n$ reactive, skill-conditioned policies that preserve diversity while satisfying the near-optimality constraint. We build on the off-policy Domino framework [11], and augment it with three key modifications inspired by efficient offline-to-online RL [36].
Following Zahavy et al. [11], we use the gradient of the diversity objective as intrinsic reward for policy optimization, that is
@@
where $\bar{\phi}_{z}=\operatorname*{\mathbb{E}}_{s\sim\rho_{\pi_{z}}}[\phi(s)]$ are the expected features per skill. For consistency, the intrinsic term measures novelty in the same projection space $\phi$ used in CNS. We balance extrinsic and intrinsic rewards as in Eq. 7 using bounded Lagrange multipliers that we update via dual ascent to enforce the near-optimality constraint. Similar to Domino, we track a running estimate of the expected features per skill, which we initialize with the population means from CNS.
To efficiently incorporate the CNS data, we use symmetric sampling [36, 37, 38], which means that each batch is composed of equal parts of online and offline transitions. Unlike prior work, we do not use the full data, however, since we found that trajectories from early iterations of CNS might fail to make meaningful interactions with the environment. So we only keep those trajectories that are nearly optimal following Eq. 1. To also exploit suboptimal data, we use a relaxed near-optimality criterion of $\alpha/4$ and estimate $v^{*}=\max_{\tau\in{\cal D}}J(\tau)$. Since this procedure rejects unequal numbers of trajectories per-skill, we adapt the symmetric sampling by balancing offline and online data equally per-skill. Further, we use a high number of learning steps per environment step (update-to-data ratio, UTD) to learn from the diverse CNS data as efficiently as possible. This allows to propagate exploration data quickly through the network, but requires extensive regularization to prevent overfitting. While multiple remedies to this issue are known, we follow Ball et al. [36] in using random ensemble distillation [39], and observation and layer normalization.
@@ 4 Related Work
Various methods to search diverse policies have been proposed. Quality-Diversity (QD) and evolutionary methods search in a gradient-free manner, populating archives of high-performing, behaviorally distinct solutions [40, 41] or co-optimizing fitness and novelty across populations [12, 10, 42]. In principle, these approaches could be used in the first step of the proposed curriculum. However, most prior work focuses on optimizing in policy parameter space before distilling policies [43, 44, 45], a less effective process as we find in this work. Recently, gradient-based RL has been reformulated to discover multiple policies via intrinsic diversity bonuses. Notably, Eysenbach et al. [26] maximize mutual information between skills and states, but focus on unsupervised skill discovery, while we target task-driven diversity in this work. For environments with extrinsic rewards, prior methods either learn policies sequentially [46, 14, 47, 48], or in parallel for greater efficiency [11, 15, 48]. Further, the methods differ in the employed optimization paradigm. While earlier works balance extrinsic and intrinsic rewards with fixed scalars [14, 49], more recent works proposed adaptive weighting schemes [12, 13]. In particular, Zahavy et al. [11] proposed constrained optimization with Lagrange multipliers, which we adopt in this paper.
@@
Exploration is a fundamental aspect of RL, enabling agents to effectively sample the environment, avoid premature convergence to suboptimal policies, and enhance both learning performance and generalization. Accordingly, numerous exploration strategies have been proposed in the literature [50]. A common strategy perturbs the agent’s actions – often via Gaussian or temporally correlated noise processes [51, 52]. Another line of work introduces parameter noise, where noise is applied directly to the agent’s parameters rather than to its actions [53, 54]. Beyond pure noise, intrinsic-reward methods augment the extrinsic task reward with bonuses for novelty. These approaches include techniques based on knowledge-based exploration, which maximizes prediction error [55], competence-based exploration [56, 26, 57, 17] and data-based exploration, which maximizes entropy [27]. While the above intrinsic exploration objectives or parameter noise explore in the parameter space, we propose to perform intrinsically motivated exploration in the much lower-dimensional trajectory space. By optimizing a novelty objective on trajectories using an ES, we can explore over entire behaviors, much like temporally correlated action noise, but with a self-optimizing noise distribution. Finally, unlike methods that rely on expert demonstrations to guide exploration [19, 58], our approach does not require prior knowledge or external supervision.
@@
Offline RL addresses the issue of data inefficiency inherent in online RL by training solely on a fixed dataset of past interactions, but it often suffers from sub-optimality due to limited or biased data [59]. To overcome these problems, the paradigm of offline-to-online RL has been proposed, where a policy is trained on offline and online data in conjunction [60, 61, 62, 63, 59, 64, 65]. The approaches broadly fall into two categories based on whether the dataset from the replay buffer is discarded after a pretraining phase, or whether it is retained [61]. Pretrain-and-discard methods first pretrain a policy or critic on the dataset and then use the same networks for finetuning without retaining any prior data [47, 66, 67]. Data retention methods keep the dataset in the replay buffer for at least a fraction of the training procedure. Many methods that fall into this category first employ an offline pre-training phase but also retain the data during online training to prevent catastrophic forgetting [60, 59, 68]. Other works instead directly train a policy online, mixing samples from the prior dataset and current rollouts [36, 69, 62, 37, 19]. We adopt this mixing strategy for our CNS dataset finetuning but observe that none of these approaches explicitly address the preservation or enhancement of policy diversity during online adaptation. In this work, we close that gap by investigating how offline–online data selection and mixing influence the retention and amplification of diverse behaviors.

### v2

@@ Abstract
Being able to solve a task in diverse ways makes agents more robust to task variations and less prone to local optima. In this context, constrained diversity optimization has emerged as a powerful reinforcement learning (RL) framework to train a diverse set of agents in parallel. However, existing constrained-diversity RL methods often under-explore in complex tasks such as robotic manipulation, leading to a lack in policy diversity. To improve diversity optimization in RL, we therefore propose a two-stage curriculum for diversity optimization. The key idea of our method is to leverage a structured spline-based trajectory prior as an inductive bias to seed diverse, high-reward behaviors before learning step-based policies. In our empirical evaluation, we provide novel insights into the shortcomings of skill-based diversity optimization, and demonstrate empirically that our curriculum improves the diversity of the learned skills.
@@ 1 Introduction
Reinforcement Learning (RL) has driven breakthroughs in robot locomotion (Hwangbo et al., 2019), game-playing (Mnih et al., 2015; Silver et al., 2017), and foundation-model finetuning (Bai et al., 2022). While effective, most RL methods assume a unimodal action distribution and produce only a single policy. In contrast, humans and animals routinely solve the same task using multiple qualitatively different strategies. Such variability is also desirable in RL, as strategy diversity increases solution quality and robustness (Page, 2017; Hong & Page, 2004). Therefore, this work considers the discovery of a policy set that maximizes the reward in diverse ways.
@@
A number of previous works have investigated this problem from various perspectives. Notably, the fields of Novelty Search (NS) and Quality-Diversity (QD) have proposed a multitude of algorithms which populate an archive of solutions based on their novelty and performance (Lehman & Stanley, 2011a; b; Conti et al., 2018). Further, gradient-based RL approaches define intrinsic diversity rewards that they combine with extrinsic task rewards using Lagrange multipliers (Zahavy et al., 2023), bandits (Parker-Holder et al., 2020), or linear combinations (Kumar et al., 2020; Masood & Doshi-Velez, 2019; Gangwani et al., 2019). While effective, the above approaches are not without shortcomings. QD may produce exceptional results, but often at the cost of sample efficiency and manual feature design. Gradient-based diversity or entropy bonuses in RL may still collapse to a few modes and remain under-evaluated in challenging contact-rich tasks (Rho et al., 2025; Emukpere et al., 2024), a finding which we corroborate in this work.
@@
Inductive biases such as hierarchical policy structures (Pateria et al., 2021), graph-based relational representations (Battaglia et al., 2018), and physics-based priors (Ramesh & Ravindran, 2023) have driven significant advances in RL. We argue that diversity optimization also benefits from inductive biases We propose a new and simple trajectory-first curriculum for learning diverse policies that explores at the level of smooth trajectories instead of neural network parameters (Section˜5). Concretely, the curriculum (i) uses an evolutionary search over open-loop action sequences to uncover a diverse set of high-reward behaviors, and (ii) distills these behaviors into distinct, off-policy, model-free policies. While prior work proposed similar formulations that first solve exploration and then learning (Campos et al., 2020; Nair et al., 2018), we do not rely on human demonstrations and propose an evolutionary approach to maximize diversity at trajectory level instead of optimizing neural-network parameters, which can be inefficient. Based on our algorithm, we empirically highlight shortcomings of existing diversity optimization methods in Section˜5 and illustrate how the proposed curriculum enables learning diverse sets of robot manipulation policies.
@@
In short, we make three contributions with this work. First, we propose a novel curriculum for diversity optimization under extrinsic task rewards (see Fig.˜1). Second, we introduce Constrained Novelty Search (CNS) to discover diverse trajectories in the first step of this curriculum (Section˜3.1). Finally, we investigate how diversity can be maintained during online training of control policies from this data (Section˜3.2).
@@
Figure 1: Overview of the proposed diversity curriculum. We use a spline-based trajectory prior to improve exploration. First, an evolution strategy explores in trajectory space to maximize novelty of trajectory parameters $\omega\in\Omega$ under performance constraints. Then, this data is used to warmstart the online training of multiple RL agents $\pi\in\Pi$ to solve the same optimization problem in policy space.
@@ 2 Preliminaries
While earlier works used scalars to balance diversity and task rewards, Zahavy et al. (2023) introduced the following constrained MDP formulation:
@@
where $\Pi^{n}$ is the current set of policies, $v^{*}\triangleq\max_{\pi\in\Pi}J(\pi)$ is the value of the optimal policy and $\alpha\in[0,1)$ is a hyperparameter controlling the optimality constraint. This constrained optimization problem can then be solved efficiently using Lagrange multipliers that are tuned using dual ascent (Altman, 2021; Borkar, 2005). Similar to subsequent work (Zheng et al., 2024), we adopt the same problem formulation in this work. To quantify diversity, we will measure the distance to the nearest neighbor, which shall be maximized:
@@
Most novelty-based approaches to skill learning maximize the entropy of the policy set by using the entropy of the current policy set as an intrinsic reward (Conti et al., 2018; Lehman & Stanley, 2011a; Liu & Abbeel, 2021). To quantify entropy, the particle-based entropy estimator by Singh et al. (2003) is commonly employed which estimates the sparsity of the distribution based on the distance between the datapoints $\{x_{i}\}_{i=1}^{n}$ and their $k$-nearest neighbor: ${\cal H}_{\text{particle}}(X)\propto\sum_{x_{i}\in X}\log\lVert x_{i}-x_{i}^{(k)}\rVert$. In practice, we choose $k=1$. In other words, we measure the particle-based entropy as distance to the nearest neighbor, and add a constant $c=1$ for numerical stability:
@@ 3 Evolutionary Exploration for Diverse Policy Discovery
[…]
@@ 3.1 Constrained Novelty Search for Spline-based Exploration
The first stage of our curriculum directly optimizes agent trajectories $\tau\in{\mathbb{R}}^{T\times u}$, where $T$ denotes the number of timesteps and $u$ the robot’s degrees of freedom. We represent a trajectory as a B-spline parameterized by a control point matrix $\omega\in{\mathbb{R}}^{m\times u}$. Following the constrained diversity optimization objective in Eq.˜1, we optimize a set of trajectory parameters $\Omega=\{\omega_{i}\}_{i=1}^{n}$ such that the resulting trajectories $\tau(\omega)$ are as diverse as possible under the constraint of near-optimality:
@@
where $J(\tau(\omega_{i}))=\textstyle\sum_{t}r_{\text{ext}}(s^{i}_{t},a^{i}_{t})$, and $v^{*}=\max_{\omega}J(\tau(\omega))$. Since the extrinsic task reward function $r_{\text{ext}}$ is generally non-differentiable we optimize this problem using a multi-population evolution strategy (ES). Following Zahavy et al. (2023) we solve the dual optimization using gradient ascent with bounded Lagrange multipliers $\{\lambda_{i}\}_{i=1}^{n}$, which yields the following trajectory reward that we evaluate on the trajectory $\tau_{i}=\tau(\omega_{i})$:
@@
where $\phi$ is a feature extraction function that projects the states to a lower dimension, $\lambda_{i}$ is the $i$-th Lagrange multiplier, and $\sigma$ denotes the sigmoid function $\sigma(x)=1/(1+\exp(-x))$. We note that Eq.˜7 is a generalized version of the novelty search objective from Conti et al. (2018), but using population-level Lagrange multipliers instead of a single heuristically selected scalar. Using Lagrange multipliers permits not only to consider different weights for different populations, but also a dynamic adaptation of these weightings. We denote this objective and its optimization as Constrained Novelty Search (CNS) in the following. While Zahavy et al. (2023) derive the analytical gradient of Eq.˜2 for their intrinsic reward, CNS approximates this gradient by stochastic sampling from an ES, which approximates natural gradient steps on the novelty objective (Akimoto et al., 2012; Glasmachers et al., 2010; Hansen & Ostermeier, 2001).
@@
Finally, we note again that we optimize in ${\mathbb{R}}^{m\times u}$ instead of ${\mathbb{R}}^{d}$ at this step where $d$ is the dimension of policy parameters. Since typically $m\times u\ll d$, we can optimize Eq.˜7 with the CMA-ES (Hansen & Ostermeier, 2001), which is empirically more sample efficient in moderately high dimensional parameter spaces than using isotropic search distributions (Salimans et al., 2017; Conti et al., 2018) .
@@ 3.2 Efficient Online Diversity Optimization from Prior Data
Given a diverse dataset ${\cal D}=\{(\tau_{i},\,z_{i},\,r_{\mathrm{ext}}(\tau_{i}))\}_{i=1}^{\ell}$ with skill labels $z_{i}\in\{1,\dots,n\}$ from CNS, our second stage learns $n$ reactive, skill-conditioned policies that preserve diversity while satisfying the near-optimality constraint. We build on the off-policy Domino framework (Zahavy et al., 2023), and augment it with three key modifications inspired by efficient offline-to-online RL (Ball et al., 2023).
Following Zahavy et al. (2023), we use the gradient of the diversity objective as intrinsic reward for policy optimization, that is
@@
where $\bar{\phi}_{z}=\operatorname*{\mathbb{E}}_{s\sim\rho_{\pi_{z}}}[\phi(s)]$ are the expected features per skill. For consistency, the intrinsic term measures novelty in the same projection space $\phi$ used in CNS. We balance extrinsic and intrinsic rewards as in Eq.˜7 using bounded Lagrange multipliers that we update via dual ascent to enforce the near-optimality constraint. Similar to Domino, we track a running estimate of the expected features per skill, which we initialize with the population means from CNS.
To efficiently incorporate the CNS data, we use symmetric sampling (Ball et al., 2023; Vecerik et al., 2017; Ross & Bagnell, 2012), which means that each batch is composed of equal parts of online and offline transitions. Unlike prior work, we do not use the full data, however, since we found that trajectories from early iterations of CNS might fail to make meaningful interactions with the environment. So we only keep those trajectories that are nearly optimal following Eq.˜1. To also exploit suboptimal data, we use a relaxed near-optimality criterion of $\alpha/4$ and estimate $v^{*}=\max_{\tau\in{\cal D}}J(\tau)$. Since this procedure rejects unequal numbers of trajectories per-skill, we adapt the symmetric sampling by balancing offline and online data equally per-skill. Further, we use a high number of learning steps per environment step (update-to-data ratio, UTD) to learn from the diverse CNS data as efficiently as possible. This allows to propagate exploration data quickly through the network, but requires extensive regularization to prevent overfitting. While multiple remedies to this issue are known, we follow Ball et al. (2023) in using random ensemble distillation (Chen et al., 2021), and observation and layer normalization.
@@ 4 Related Work
Various methods to search diverse policies have been proposed. Quality-Diversity (QD) and evolutionary methods search in a gradient-free manner, populating archives of high-performing, behaviorally distinct solutions (Mouret & Clune, 2015; Cully et al., 2015) or co-optimizing fitness and novelty across populations (Parker-Holder et al., 2020; Conti et al., 2018; Vassiliades et al., 2017; Braun et al., 2025). In principle, these approaches could be used in the first step of the proposed curriculum. However, most prior work focuses on optimizing in policy parameter space before distilling policies (Faldor et al., 2023; Macé et al., 2023; Chalumeau et al., 2023), a less effective process as we find in this work. Recently, gradient-based RL has been reformulated to discover multiple policies via intrinsic diversity bonuses. Notably, Eysenbach et al. (2019) maximize mutual information between skills and states, but focus on unsupervised skill discovery, while we target task-driven diversity in this work. For environments with extrinsic rewards, prior methods either learn policies sequentially (Fu et al., 2023; Masood & Doshi-Velez, 2019; Zhou et al., 2022; Chen et al., 2024), or in parallel for greater efficiency (Zahavy et al., 2023; Gangwani et al., 2019; Chen et al., 2024; Celik et al., 2024). Further, the methods differ in the employed optimization paradigm. While earlier works balance extrinsic and intrinsic rewards with fixed scalars (Masood & Doshi-Velez, 2019; Liu et al., 2017), more recent works proposed adaptive weighting schemes (Parker-Holder et al., 2020; Kumar et al., 2020). In particular, Zahavy et al. (2023) proposed constrained optimization with Lagrange multipliers, which we adopt in this paper.
@@
Exploration is a fundamental aspect of RL, enabling agents to effectively sample the environment, avoid premature convergence to suboptimal policies, and enhance both learning performance and generalization. Accordingly, numerous exploration strategies have been proposed in the literature (Ladosz et al., 2022). A common strategy perturbs the agent’s actions – often via Gaussian or temporally correlated noise processes (Fujimoto et al., 2018; Hollenstein et al., 2022). Another line of work introduces parameter noise, where noise is applied directly to the agent’s parameters rather than to its actions (Plappert et al., 2018; Fortunato et al., 2018). Beyond pure noise, intrinsic-reward methods augment the extrinsic task reward with bonuses for novelty. These approaches include techniques based on knowledge-based exploration, which maximizes prediction error (Burda et al., 2019), competence-based exploration (Houthooft et al., 2016; Eysenbach et al., 2019; Laskin et al., 2022; Zheng et al., 2024) and data-based exploration, which maximizes entropy (Liu & Abbeel, 2021). While the above intrinsic exploration objectives or parameter noise explore in the parameter space, we propose to perform intrinsically motivated exploration in the much lower-dimensional trajectory space. By optimizing a novelty objective on trajectories using an ES, we can explore over entire behaviors, much like temporally correlated action noise, but with a self-optimizing noise distribution. While prior work investigated RL at trajectory level (Otto et al., 2023; Klink et al., 2020; Celik et al., 2024) to improve exploration, these works do not learn step-based reactive policies but predict full action sequences, which is a drawback in practical domains such as robotics. Finally, unlike methods that rely on expert demonstrations to guide exploration (Nair et al., 2018; Salimans & Chen, 2018), our approach does not require prior knowledge or external supervision.
@@
Offline RL addresses the issue of data inefficiency inherent in online RL by training solely on a fixed dataset of past interactions, but it often suffers from sub-optimality due to limited or biased data (Liu et al., 2024). To overcome these problems, the paradigm of offline-to-online RL has been proposed, where a policy is trained on offline and online data in conjunction (Fang et al., 2022; Nakamoto et al., 2023; Zhou et al., 2025; Nakamoto et al., 2023; Feng et al., 2024; Liu et al., 2024; Wang et al., 2023). The approaches broadly fall into two categories based on whether the dataset from the replay buffer is discarded after a pretraining phase, or whether it is retained (Zhou et al., 2025). Pretrain-and-discard methods first pretrain a policy or critic on the dataset and then use the same networks for finetuning without retaining any prior data (Zhou et al., 2022; Uchendu et al., 2023; Wolczyk et al., 2024). Data retention methods keep the dataset in the replay buffer for at least a fraction of the training procedure. Many methods that fall into this category first employ an offline pre-training phase but also retain the data during online training to prevent catastrophic forgetting (Fang et al., 2022; Liu et al., 2024; Hu et al., 2023). Other works instead directly train a policy online, mixing samples from the prior dataset and current rollouts (Ball et al., 2023; Song et al., 2023; Nakamoto et al., 2023; Vecerik et al., 2017; Nair et al., 2018). We adopt this mixing strategy for our CNS dataset finetuning but observe that none of these approaches explicitly address the preservation or enhancement of policy diversity during online adaptation. In this work, we close that gap by investigating how offline–online data selection and mixing influence the retention and amplification of diverse behaviors.